Write down expectation value of r for hydrogen atom.
1
Expert's answer
2020-11-05T07:35:45-0500

Expectation value of r for hydrogen atom:


"\big<r\big>=\int^\infty_0\psi_{1s}(r)r\psi_{1s}(r)4\pi r^2 \text{d}r=\frac{3}{2}"

Or


"\big<r\big>=\frac32 a_0."